### v1

# Impact of Nitrogen and Oxygen Interstitials on Niobium SRF Cavity Performance

Abstract
Superconducting radio-frequency (SRF) cavities are the leading technology for highly efficient particle acceleration, and their performance can be significantly enhanced through the controlled introduction of interstitial impurities into bulk niobium. Nitrogen doping has demonstrated a substantial reduction in surface resistance losses, which improves the quality factor of the cavities. More recently, oxygen doping has emerged as a promising alternative, demonstrating comparable reductions in surface resistance. In this study, we combine cavity measurements on niobium SRF cavities subjected to a range of nitrogen- and oxygen-based treatments with material characterizations performed on cavity cutouts processed under identical conditions. This approach allows us to quantitatively assess the contribution of each impurity to the reduction of surface resistance. We find that nitrogen is ten times more effective than oxygen in reducing surface resistance at . We propose a model to explain this variation, suggesting that nitrogen more effectively traps hydrogen, thus suppressing the formation of niobium hydrides within the RF penetration layer and enabling an improved superconducting gap.
Extending the performance of superconducting radiofrequency (SRF) cavities is critical for enabling next-generation superconducting accelerators, with higher energies, reduced power losses, and lower cryogenic costsPadamsee, Knobloch, and Hays (2008); Isagawa (1980). In SRF cavities, electromagnetic fields build up within the cells; these fields accelerate charged particles along the cavity axis and, at the same time, induce surface currents that penetrate into the superconducting surface, defining the RF layerIsagawa (1980); Maxfield and McLean (1965). Over the past decade, significant improvements in performance have been made by modifying the composition of the RF layer through the introduction of interstitial impurities.
Surface treatments such as nitrogen doping, nitrogen infusion, and more recently, oxygen-based baking treatments, have been developed to enhance SRF performance. Nitrogen doping is one surface treatment which introduces uniform, dilute concentrations of N throughout the RF layer and yields high quality factors (Q0) Grassellino et al. (2013). In contrast, nitrogen infusion is a treatment which deposits a higher concentration of N confined to the first of the surface, enabling higher accelerating gradients (Eacc) rather than ultra-high Q0Grassellino et al. (2018). While N doping has been successfully industrialized and implemented in full-scale cryomodules, N infusion has proven more challenging to translate and apply outside of Fermilab Galayda (2018); Raubenheimer (2018). More recently, oxygen-based surface treatments have recently emerged as a promising and reliable alternative Bafia, Grassellino, and Romanenko (2022); Ito et al. (2021); Lechner et al. (2021). These treatments require only a single baking step with no subsequent electropolishing (EP), taking advantage of the naturally occurring Nb oxide; baking dissolves the oxide and drives inward O diffusion Bafia, Grassellino, and Romanenko (2022); Veit et al. (2019); Bafia et al. (2024). Oxygen treatments can be categorized by temperature. Low-temperature baking at enhances accelerating gradients in a manner similar to N infusion, with O concentrated in the first Bafia et al. (2019); Bafia, Grassellino, and Romanenko (2022); Padamsee (2009). Medium-temperature (mid-T) baking () yields more uniform distributions of O and increases Q0 akin to N doping Posen et al. (2020); Ito et al. (2021). These treatments have also been applied to full-scale cryomodules and are under consideration for future major accelerators, including two-step low temperature baking for the International Linear Collider and mid-T baking for the Future Circular ColliderKatayama et al. (2022); Bafia et al. (2019); Pan et al. (2024). Collectively, these results demonstrate that nitrogen- and oxygen-based approaches can be tailored for either high-gradient or high-Q0 performance, motivating a direct comparison of their underlying mechanisms and the magnitudes of their effects.
| Cavity | Treatment Steps | |
| TE1AES010 | EP Baseline | |
| TE1AES010 | 120 ∘C 3 h in-situ | + 120 ∘C 3 h (total 6 h) in-situ |
| TE1PAV009 | 120 ∘C 48 h in-situ | |
| TE1AES017 | 200 ∘C 1 h in-situ | + 200 ∘C 10 h (total 11 h) in-situ |
| TE1AES021 | 200 ∘C 20 h in-situ | |
| TE1AES024 | 2/0+5 µm N Doped | |
| TE1RI003 | 3/60+10 µm N Doped | + 8 µm EP (3/60+18 µm N Doped) |
| TE1PAV012 | N Infused (120 ∘C 48 h with 25 mTorr N2 injection) | |
The performance of SRF cavities is limited by the RF surface resistance, which can be expressed as the sum of a temperature-dependent resistance (RT) and a temperature-independent residual resistance (R. RT arises from dissipation associated with thermally excited quasiparticles in the superconducting state and can be described using Mattis-Bardeen theory, an extension of BCS theory which accounts for an applied RF field Mattis and Bardeen (1958); Bardeen, Cooper, and Schrieffer (1957). Interstitial impurities such as nitrogen and oxygen influence RT by modifying the mean free path in the RF layer, which can optimize the superfluid density and suppress quasiparticle losses. RT can be expressed as: where is a parameter dependent on the electronic mean free path, is the cavity resonant frequency, is the superconducting energy gap and is Boltzmann’s constant Halbritter (1974). Rres can be caused by an increase in the number of subgap quasiparticle states, lossy oxides, and metallic inclusions, and vortex oscillations Kubo and Gurevich (2019).
Despite empirical evidence for the benefits of N and O interstitials in the RF layer, the underlying physical mechanisms are not yet fully understood. Theoretical models suggest that interstitials modify key superconducting parameters, including the mean free path, penetration depth, and superconducting gap Gurevich (2017, 2014). A mismatch between a doped (dirty) surface and a clean bulker can increase the superheating field and delay vortex penetration under increasing RF fieldsKubo and Gurevich (2019); Gurevich (2015). In the nonlinear Meissner screening theory, interstitials optimize the mean free path, increasing the superfluid density and reducing losses Sauls (2022). At higher RF fields, nonlinear effects, including pair breaking and vortex penetration, can increase RT, contributing to field-dependent degradation of cavity performance Sauls (2022). Nonequilibrium superconductivity is another mechanism in which interstitial impurities modify the quasiparticle distribution from the thermal equilibrium distribution by enhancing recombination or altering inelastic scattering rates under RF excitationMartinello et al. (2018). A microscopic description of these nonequilibrium effects is provided by Eliashberg theory, in which interstitial impurities modify gap anisotropy and inelastic scattering under the strong-coupling regime Eliashberg (1970).
Another proposed mechanism for enhancing SRF cavity performance is that interstitial impurities act as hydrogen traps Ford, Cooley, and Seidman (2013a); Isagawa (1980). Hydrogen is highly diffusive in Nb, making it particularly susceptible to hydrogen uptake. Additionally, many standard cavity processing techniques rely on acids, which may introduce additional H Knobloch (2003); Barkov et al. (2013). When Nb is cooled below , interstitial H may precipitate as niobium nanohydrides, which become superconducting through the proximity effect, thus degrading superconducting properties and increasing losses Romanenko et al. (2013a); Knobloch (2003). The formation of these nanohydrides has been observed with cryogenic transmission electron microscopy and cryogenic atomic force microscopyTrenikhina et al. (2015); Sung et al. (2024). Introducing dilute concentration of impurities, such as N, O, C, and Ti, can distort the bcc Nb lattice and allow the interstitials to act as effective traps for HKoufalis et al. (2017); Ford, Cooley, and Seidman (2013b); Dhakal et al. (2013). First-principle calculations confirm it is more energetically favorable for interstitial H to bind to N or O than to NbFord, Cooley, and Seidman (2013b).
In this work, we conduct a quantitative analysis of the effect of nitrogen and oxygen impurities on RT by correlating cavity measurements to materials characterizations from time-of-flight secondary ion mass spectropy (ToF-SIMS) and atom probe tomography (APT). Our results indicate that nitrogen is up to ten times more effective than oxygen at at reducing RT. We interpret these findings in the context of previous experimental studies, Mattis-Bardeen theory, and theoretical models, and we propose a framework that connects improvements in cavity performance to changes in underlying superconducting material properties.
Single-cell TESLA-shaped cavities fabricated from high-purity niobium were used to investigate the effects of N and O interstitials on RF performance. Cavity cutouts, in diameter, from TE1AES008 were prepared using the same surface treatment procedures as the cavities. Initial surface preparation followed a standard baseline procedure of a bulk 120 µm EP removal, followed by degassing for 3 h in an ultra-high vacuum (UHV) furnace, 40 µm EP removal, and high-pressure rinsing to clean the surface Padamsee, Knobloch, and Hays (2008). Subsequently, the treatments detailed in Table 1 were applied. In-situ baked cavities were assembled prior to baking and maintained vacuum through testing, whereas N doped, N infused, and EP cavities were assembled post-treatment. After assembly, the vacuum level within each cavity was maintained below Torr.
To quantify the concentration of N and O impurities introduced into the cutouts after each processing technique outlined in Table 1, we employed two complementary methods. First, we used ToF-SIMS to obtain depth profiles of the impurities present in the Nb lattice. Implanted N and O standards were measured in parallel to calibrate the relative ToF-SIMS intensity to absolute concentration, allowing a direct comparision between O and N impuritiesBudrevich and Hunter (1998). Specifics on ToF-SIMS measurement are described in the supplementary material. Then, as an independent measure of concentration, we analyzed one sample from CAV018, treated with a 2/0+5µm N doping recipe with APT. Nanotips were prepared following the procedures described in Lee et al. (2024). APT measurement details are provided in the supplementary material. The same sample from CAV018 was also measured with ToF-SIMS to validate the absolute O and N concentrations between the two measurement techniques.
RF performance of each single-cell Nb cavity was evaluated at the Fermilab Vertical Test Stand (VTS) in continuous-wave operation using power balance measurementsMelnychuk, Grassellino, and Romanenko (2014). All cavities underwent a fast cooldown protocol to minimize magnetic flux trappingRomanenko et al. (2014); Posen et al. (2016). Q0 vs. Eacc curves were measured at both and to decompose the surface resistance into RT and RresMartinello et al. (2016). The temperature-dependent component at can be obtained from RGQ0() GQ0(), where G = 270 is a geometry factor that is only dependent on cavity shape. This study primarily investigates the evolution in RT as our main interest is the effect of bulk impurities on superconducting properties.
Figure 1 shows ToF-SIMS depth profile of O and N in Nb cavity cutouts for each treatment. The EP baseline exhibits a clean surface with no detectable O or N impurities. O-based treatments introduce significant concentrations of O. Baking at gradually dissolves oxygen into the bulk with a diffusion length of , and baking at leads to a more uniform oxygen distribution within the RF layer Bafia, Grassellino, and Romanenko (2022). No N was detected in the EP, 120∘C, or 200∘C baked samples above background levels. N doped samples show uniform and dilute concentrations of 103 ppma of N extending well beyond the RF layer, with negligible O concentration. In contrast, the N infused sample contains both N is in the first 20 nm and O comparable to a 120∘C48 h bake. This enables consideration of treatments that introduce both O and N.
To validate the concentrations from ToF-SIMS, APT was performed on a N Doped sample (Fig. 2). We identified a cylindrical region of interest (ROI) to be representative of the average surface (see supplementary material further discussion). Within the ROI, the average N concentration was ppma. The same sample later measured with ToF-SIMS yielded average N concentration of ppma, corresponding to a Z-score of 0.43. A direct comparison of the bulk O concentration from a depth of gives ppma for APT and ppma for ToF-SIMS, corresponding to a Z-score of 0.30. The agreement in both N and bulk O concentrations across independent techniques provides strong confirmation of the reliability of the measurements.
Figure 3 plots the corresponding RF measurements of RT as a function of Eacc for each treatment. Longer 120∘C bakes yield a gradual increase in maximum Eacc, consistent with previous studies Romanenko et al. (2013b); Bafia, Grassellino, and Romanenko (2022). Both 200∘C in-situ baking and N doping substantially reduce RT. Notably, the performance of the 200∘C20 h baked cavity is comparable to that of both the 2/0+5 µm and 3/60+10 µm N doped treatments, indicating that O and N can play analogous roles in mitigating BCS-related losses.
A direct correlation between cavity RF performance and impurity concentration is shown in Fig. 4. We extracted RT at 16 MV/m for each of the treatments from Fig. 3. 11116 MV/m was chosen as it avoids both high field Q-slope degradation at 25 MV/m and low field (5 MV/m) effects from surface oxides. and averaged impurity concentrations from Fig. 1 over the RF layer, excluding the top to remove surface oxide contributions. RT for both O- and N-based treatments exhibit approximately logarithmic scaling with concentration, where higher interstitial concentrations systematically lower RT. The lines of best fit indicate that nitrogen provides a stronger suppression per unit concentration than oxygen. At 16 MV/m, N doped cavities display RT values similar to those of 200∘C20 h but with an order of magnitude less interstitial content. Thus, either species can drive comparable performance improvements when present at sufficient levels. It is somewhat surprising that there is such a large difference in the effects of N and O, given that both N and O can only trap one H per atom Morkel, Wipf, and Neumaier (1978); Richter and Springer (1978).
N infusion provides a hybrid case. In Fig. 4, it is represented twice: once by its O concentration, comparable to a 120∘C48 h bake, and once by its N concentration. Relative to a 120∘C48 h bake, the N infused cavity shows an additional reduction in RT of 2.5 n, attributable to the small addition of N. This suggests that the combined presence of O and N interstitials can have an additive effect in lowering RT.
To contextualize these results, we compare them with BCS theory and previous experimental studies. Mean free path analysis (Fig. 5) provides a unifying framework for O and N. Average impurity concentrations were converted into mean free path () using , where is the absolute impurity concentration and the atomic radius of Nb. For baked samples, corresponds to average O concentration; for N doped and N infused samples, is the sum of O and N concentrations. For reference, the theoretical BCS curves for various were calculated using SRIMP with fixed parameters: K, GHz, nm, and nm Halbritter (1970). Our results are also compared with experimentally determined values from Martinello et al. (2016) and Bafia (2020), obtained through low-energy muon spin rotation measurements (LE-µSR) and cavity fits. We observe good agreement between extracted for EP and N doped cavities across four independent measurement techniques: ToF-SIMS, APT, LE-µSR, and cavity fits with RF measurements.
Overall, the data highlight two distinct regimes. Treatments that confine impurities to the near-surface, such as baking at and N infusion, primarily reduce without significantly altering the superconducting gap. This results in significant improvement in Eacc but only limited improvement in Q0. It is likely that these treatments form a dirty—clean superconducting bilayer, which can enhance the superheating field beyond its theoretical limits by pushing superconducting currents deeper into the clean bulk Gurevich (2017); Kubo (2016). In contrast, N doping reduces and simultaneously increases , resulting high Q0 performance. Similarly, longer durations of baking at progressively lower RT by enhancing , even when the average impurity concentration changes minimally. These results demonstrate that both O and N reduce RT through a combination of mean free path optimization and gap enhancement, with gap enhancement observed only in treatments exhibiting sufficient uniformity in the impurity depth profile throughout the RF layer. A more uniform impurity profile likely ensures that the entire RF penetration layer remains within the optimal scattering regime and enables a more homogeneous superconducting state.
Supporting this, N doped cavities exhibit both a larger average and a lower inelastic scattering parameter () compared to EP cavities, indicative of an improved quasiparticle condensate Bafia et al. (2025). Furthermore, point-contact tunneling spectroscopy (PCTS) studies have demonstrated that cavities treated at for have inelastic scattering rates nearly twice as large as those in N doped samplesGroll et al. (2018). This indicates that a homogeneous interstitial layer suppresses , thereby narrowing the quasiparticle density of states and reducing quasiparticle losses. These improvements in the condensate structure and electronic homogeneity are directly linked to enhanced BCS behavior, as discussed in recent theoretical work emphasizing the role of spatial disorder and scattering on the electrodynamics of superconductors under RF fields Kubo (2022); Herman and Hlubina (2021); Zarea, Ueki, and Sauls (2023).
For nitrogen, the increase in the average superconducting gap and the narrowing of its distribution is achieved with approximately ten times lower concentration. This suggests that nitrogen has a stronger effect per atom, likely due to more effective hydrogen trapping. First-principle calculations by Ford, Cooley, and Seidman (2013b) support this, indicating comparable binding energies for H-O (7.02 eV) and H-N (7.39 eV), though this is not sufficient to fully explain the tenfold difference. One possible explanation is that N doped surfaces form fewer vacancies and less defective suboxides than oxygen-doped surfaces, thereby enhancing their superconducting properties Yang et al. (2018); Wenskat et al. (2022). X-ray photoelectron spectroscopy and transmission electron microscopy studies have also indicated that N doped surfaces form a more stoichmetric Nb2O5 layer with thinner metallic suboxide layers Lechner et al. (2020); Fang et al. (2023).
In this work, we systematically investigated the effects of O and N impurities on the performance of niobium SRF cavities and their superconducting properties. By correlating material characterizations with RF performance, we demonstrated that O and N independently modify the superconducting properties of the RF layer. N doping achieves comparable RT reduction at lower impurity concentration than O-based treatments. The reduction in RT is likely attributable gap enhancement and homogenization, as the presence of impurities may eliminate defects in the Nb lattice. Despite its lower concentration, nitrogen likely has a stronger effect due to its surface oxide having fewer defects and a higher hydrogen trapping efficiency. In comparison, the effect of oxygen is more gradual as it must overcome losses introduced by increased the vacancies in the surface oxide. This gradual effect, coupled with the simpler processing requirements for O-based treatments, highlights the potential of oxygen diffusion for finer tuning of the RF surface in future treatments. Additionally, the additive effect of O and N impurities in reducing RT for N infusion indicates potential for combined O and N impurity tailoring. Further theoretical and experimental work is needed to fully understand the distinct roles of oxygen and nitrogen at the atomic scale. Future considerations include extending the first-principle calculations to H trapping of an interstitial O or N in Nb and examining the differences in magnitude of the tunneling barrier in Nb between different vacancies in the presence of either N or O impuritiesFord, Cooley, and Seidman (2013b); Abogoda et al. (2025).
ACKNOWLEDGEMENTS
The authors would like to thank Alexandr Netepenko, Adam Clairmont, Dieter Isheim, Akshay Murthy, Damon Bice, Tim Ring, and Davida Smith for their support in sample and cavity preparation, data acquisition, and data analysis. This manuscript has been authored by the FermiForward Discovery Group, LLC under contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work was supported by the University of Chicago. Atom probe tomography was performed with the LEAP5000XS tomograph at the Northwestern University Center for Atom-Probe Tomography (NUCAPT, RRID: SCR_017770), acquired and upgraded with equipment grants from the NSF-MRI (DMR-0420532), DURIP-ONR (N00014-0400798, N00014-0610539, N00014-0910781, N00014-1712870, N00014-2312593) and DURIP-ARO (W911NF2110052) programs. NUCAPT received support from the MRSEC program (NSF DMR-2308691) at the Materials Research Center, the SHyNE Resource (NSF ECCS-2025633), and the Paula M. Trienens Institute for Sustainability and Energy at Northwestern University.