### v2

# Impact of Nitrogen and Oxygen Interstitials on Niobium SRF Cavity Performance

Abstract
Superconducting radiofrequency (SRF) cavities are the leading technology for highly efficient particle acceleration, and their performance can be significantly enhanced through the controlled introduction of interstitial impurities into bulk niobium. Nitrogen doping has demonstrated a substantial reduction in surface resistance, which improves the quality factor of the cavities. More recently, oxygen doping has emerged as a promising alternative, demonstrating comparable reductions in surface resistance. In this study, we combine cavity measurements on niobium SRF cavities subjected to a range of nitrogen- and oxygen-based treatments with material characterizations performed on cavity cutouts processed under identical conditions. This approach allows us to quantitatively assess the contribution of each impurity to the reduction of surface resistance. We find that nitrogen is ten times more effective than oxygen in reducing surface resistance at . We also observe an additive effect of O and N impurities in reducing RT. We discuss these results in the context of field dependent nonequilibrium superconductivity, gap enhancement, and hydrogen trapping mechanisms.
Extending the performance of superconducting radiofrequency (SRF) cavities is critical for enabling next-generation superconducting accelerators, with higher energies, reduced power losses, and lower cryogenic costsPadamsee et al. (2008); Isagawa (1980). In SRF cavities, electromagnetic fields build up within the cells; these fields accelerate charged particles along the cavity axis and, at the same time, induce surface currents that penetrate into the superconducting surface, defining the RF layerIsagawa (1980); Maxfield and McLean (1965). Over the past decade, significant improvements in performance have been made by modifying the composition of the RF layer through the introduction of interstitial impurities.
Surface treatments such as nitrogen doping, nitrogen infusion, and more recently, oxygen-based baking treatments, have been developed to enhance SRF performance. Nitrogen doping is one surface treatment which introduces uniform, dilute concentrations of N throughout the RF layer and yields high quality factors (Q0) Grassellino et al. (2013). In contrast, nitrogen infusion is a treatment which deposits a higher concentration of N confined to the first of the surface, enabling higher accelerating gradients (Eacc) rather than ultra-high Q0Grassellino et al. (2018). While N doping has been successfully industrialized and implemented in full-scale cryomodules, N infusion has proven more challenging to translate and apply outside of Fermilab Galayda (2018); Raubenheimer (2018); Grassellino et al. (2017). More recently, oxygen-based surface treatments have recently emerged as a promising and reliable alternative Bafia et al. (2022); Ito et al. (2021); Lechner et al. (2021). These treatments require only a single baking step with no subsequent electropolishing (EP), taking advantage of the naturally occurring Nb oxide; baking dissolves the oxide and drives inward O diffusion Bafia et al. (2022); Veit et al. (2019); Bafia et al. (2024). Oxygen treatments can be categorized by temperature. Low-temperature baking at enhances accelerating gradients in a manner similar to N infusion, with O concentrated in the first Bafia et al. (2019, 2022); Padamsee (2009). Medium-temperature (mid-T) baking () yields more uniform distributions of O and increases Q0 akin to N doping Posen et al. (2020); Ito et al. (2021). These treatments have also been applied to full-scale cryomodules and are under consideration for future major accelerators, including two-step low temperature baking for the International Linear Collider and mid-T baking for the Future Circular Collider and Shanghai High Repetition Rate XFEL and Extreme Light Facility (SHINE) Katayama et al. (2022); Bafia et al. (2019); Pan et al. (2024). Collectively, these results demonstrate that nitrogen- and oxygen-based approaches can be tailored for either high-gradient or high-Q0 performance, motivating a direct comparison of their underlying mechanisms and the magnitudes of their effects.
| Cavity | Treatment Steps | |
| TE1AES010 | EP Baseline | |
| TE1AES010 | 120 ∘C 3 h in-situ | + 120 ∘C 3 h (total 6 h) in-situ |
| TE1PAV009 | 120 ∘C 48 h in-situ | |
| TE1AES017 | 200 ∘C 1 h in-situ | + 200 ∘C 10 h (total 11 h) in-situ |
| TE1AES021 | 200 ∘C 20 h in-situ | |
| TE1AES024 | 2/0+5 µm N Doped | |
| TE1RI003 | 3/60+10 µm N Doped | + 8 µm EP (3/60+18 µm N Doped) |
| TE1PAV012 | N Infused (120 ∘C 48 h with 25 mTorr N2 injection) | |
| TE1PAV008 | 350 ∘C 2.5 h in-situ | oxidized in air |
The performance of SRF cavities is limited by the RF surface resistance, which can be expressed as the sum of a temperature-dependent resistance (RT) and a temperature-independent residual resistance (R. RT arises from dissipation associated with thermally excited quasiparticles in the superconducting state and can be described using Mattis-Bardeen theory, an extension of BCS theory which accounts for an applied RF field Mattis and Bardeen (1958); Bardeen et al. (1957). Interstitial impurities such as nitrogen and oxygen influence RT by modifying the mean free path in the RF layer, which can optimize the superfluid density and suppress quasiparticle losses. RT can be expressed as: where is a parameter dependent on the electronic mean free path, is the cavity resonant frequency, is the superconducting energy gap and is Boltzmann’s constant Halbritter (1974). Rres can arise from an increase in the number of subgap quasiparticle states, lossy oxides, and metallic inclusions, and vortex oscillations Kubo and Gurevich (2019); Bafia et al. (2025a).
Despite empirical evidence for the benefits of N and O interstitials in the RF layer, the underlying physical mechanisms are not yet fully understood. Theoretical models suggest that interstitials modify key superconducting parameters, including the mean free path, penetration depth, and superconducting gap Gurevich (2017, 2014). A mismatch between a doped (dirty) surface and a clean bulk can increase the superheating field and delay vortex penetration under increasing RF fieldsNgampruetikorn and Sauls (2019); Kubo and Gurevich (2019); Gurevich (2015). In the nonlinear Meissner screening theory, interstitials optimize the mean free path, increasing the superfluid density and reducing losses Sauls (2022). At higher RF fields, nonlinear effects, including pair breaking and vortex penetration, can increase RT, contributing to field-dependent degradation of cavity performance Sauls (2022). Nonequilibrium superconductivity is another mechanism in which interstitial impurities modify the quasiparticle distribution from thermal equilibrium by enhancing recombination or altering inelastic scattering rates under RF excitationMartinello et al. (2018); Gurevich (2014). Eliashberg theory provides a microscopic description of these nonequilibrium effects, in which interstitial impurities modify gap anisotropy and inelastic scattering under the strong-coupling regime Eliashberg (1970).
Another proposed mechanism for enhancing SRF cavity performance is that interstitial impurities act as hydrogen traps Ford et al. (2013b); Isagawa (1980); Pfeiffer and Wipf (1976). Hydrogen is highly diffusive in Nb, making it particularly susceptible to hydrogen uptake. Additionally, many standard cavity processing techniques rely on acids, which may introduce additional H Knobloch (2003); Barkov et al. (2013). When Nb is cooled below , interstitial H may precipitate as niobium nanohydrides, which become superconducting through the proximity effect, thus degrading superconducting properties and increasing losses Romanenko et al. (2013a); Knobloch (2003). The formation of these nanohydrides has been observed with cryogenic transmission electron microscopy and cryogenic atomic force microscopyTrenikhina et al. (2015); Sung et al. (2024). Introducing dilute concentrations of impurities, such as N, O, C, and Ti, can distort the bcc Nb lattice and allow the interstitials to act as effective traps for HKoufalis et al. (2017); Ford et al. (2013a); Dhakal et al. (2013). First-principle calculations confirm it is more energetically favorable for interstitial H to bind to N or O than to NbFord et al. (2013a).
In this work, we conduct a quantitative analysis of the effect of nitrogen and oxygen impurities on RT by correlating cavity measurements to materials characterizations from time-of-flight secondary ion mass spectroscopy (ToF-SIMS). Our results indicate that nitrogen is up to ten times more effective than oxygen at at reducing RT. We interpret these findings in the context of previous experimental studies, Mattis-Bardeen theory, and theoretical models.
Single-cell TESLA-shaped cavities fabricated from high-purity niobium were used to investigate the effects of N and O interstitials on RF performance. Cavity cutouts, in diameter, from TE1AES008 were prepared using the same surface treatment procedures as the cavities. Initial surface preparation followed a standard baseline procedure of a bulk 120 µm EP removal, followed by degassing for 3 h in an ultra-high vacuum (UHV) furnace, 40 µm EP removal, and high-pressure rinsing to clean the surface Padamsee et al. (2008). Subsequently, the treatments detailed in Table 1 were applied. All but one of the in-situ baked cavities were assembled prior to baking and maintained vacuum through testing, whereas N doped, N infused, and EP cavities were assembled post-treatment. TE1PAV008 (data from Posen et al. (2020)) was in-situ baked, vented with nitrogen, disassembled, exposed to cleanroom air for 10 min, and reassembled. After assembly, the vacuum level within each cavity was maintained below Torr.
To quantify the concentration of N and O impurities introduced into the cutouts after each processing technique outlined in Table 1, we used ToF-SIMS to obtain depth profiles of the impurities present in the Nb lattice. Implanted N and O standards were measured in parallel to calibrate the relative ToF-SIMS intensity to absolute concentration, allowing a direct comparison between O and N impuritiesBudrevich and Hunter (1998). Specifics on ToF-SIMS measurement are described in the supplementary material.
RF performance of each single-cell Nb cavity was evaluated at the Fermilab Vertical Test Stand (VTS) in continuous-wave operation using power balance measurementsMelnychuk et al. (2014). All cavities underwent a fast cooldown protocol to minimize magnetic flux trappingRomanenko et al. (2014); Posen et al. (2016). Q0 vs. Eacc curves were measured at both and to decompose the surface resistance into RT and RresMartinello et al. (2016). The temperature-dependent component at can be obtained from RGQ0() GQ0(), where G = 270 is a geometry factor that is only dependent on cavity shape. This study primarily investigates the evolution in RT as our main interest is the effect of bulk impurities on superconducting properties.
Figure 1 shows ToF-SIMS depth profile of O and N in Nb cavity cutouts for each treatment. Concentrations are presented in the dimensionless unit of parts per million atomic (ppma), which represents the number of impurity atoms per total atoms. The EP baseline exhibits a clean surface with no detectable O or N impurities. Baking at gradually dissolves oxygen into the bulk with a diffusion length of , and baking at leads to a more uniform oxygen distribution within the RF layer Bafia et al. (2022). No N was detected in the EP, 120∘C, or 200∘C baked samples above background levels. N doped samples show uniform and dilute concentrations of 103 ppma of N extending well beyond the RF layer, with negligible O concentration. The N infused sample contains N in the first 20 nm and O comparable to that of a 120∘C48 h bake. A mid-T baked (350∘C2.5 h) sample displays N in the first 20 nm and a uniform concentration of O. For both N infused and mid-T baked samples, we detected nitrogen within the bulk above background levels. These treatments enable consideration of treatments that introduce both O and N. It is unclear what the source of N is in mid-T baking, but these findings mirror the nitrogen interstitials observed in Posen et al. (2020) of mid-T baked samples. Additional ToF-SIMS depth profiles of C, OH, H, and NbH are provided in the supplementary material along with quantification of trace contaminants that are only present in the first 5 nm of the surface.
Figure 2 plots the corresponding RF measurements of RT as a function of Eacc for each treatment. The reported uncertainty is 10%, an upper limit on the error Melnychuk et al. (2014). Longer 120∘C bakes yield a gradual increase in maximum Eacc, consistent with previous studies Romanenko et al. (2013b); Bafia et al. (2022). Both 200∘C in-situ baking and N doping substantially reduce RT. Notably, the performances of the 200∘C20 h baked and 350∘C2.5 h baked cavities are comparable to that of both the 2/0+5 µm and 3/60+10 µm N doped treatments, indicating that O and N can play analogous roles in mitigating BCS-related losses. No significant change in RT is expected if instead of in-situ baking, the cavities were furnace baked and re-exposured to air to allow for oxide formation. The supplementary material presents an additional study in which a partially dissolved oxide is fully regrown with no changes in RT.
A direct correlation between cavity RF performance and impurity concentration is shown in Fig. 3. We extracted RT at 16 MV/m for each of the treatments from Fig. 2 and averaged impurity concentrations from Fig. 1 over the RF layer, excluding the top to remove surface oxide contributions. RT for both O- and N-based treatments exhibit approximately logarithmic scaling with concentration, where higher interstitial concentrations systematically lower RT. The lines of best fit indicate that nitrogen provides a stronger suppression per unit concentration than oxygen. At 16 MV/m, N doped cavities display RT values similar to those of 200∘C20 h but with an order of magnitude less interstitial content. Thus, either species can drive comparable performance improvements when present at sufficient levels. It is surprising that there is such a large difference in the effects of N and O, given that both N and O can only trap one H per atom Morkel et al. (1978); Richter and Springer (1978). It may be that hydrogen trapping is not the sole mechanism which drives differences between N and O.
N infusion and mid-T baking provide hybrid cases. In Fig. 3, N infusion represented twice: once by its O concentration, comparable to a 120∘C48 h bake, and once by its N concentration. Relative to a 120∘C48 h bake, the N infused cavity shows an additional reduction in RT of 2.5 n, attributable to the small addition of N. This suggests that the combined presence of O and N interstitials can have an additive effect in lowering RT. A similar additive effect can also be observed for mid-T baking.
To contextualize these results, we compare them with BCS theory and previous experimental studies. Mean free path analysis (Fig. 4) provides a unifying framework for O and N. To take into consideration the evolution of RT with field, we extracted RT at both low field (5 MV/m) and medium field (16 MV/m). The supplemental material contains analysis at high fields (25 MV/m). Average impurity concentrations were converted into mean free path () using , where is the ion concentration in ions/cm3and Å is the atomic radius of Nb. Ion concentration can be obtained from ppma using the Nb atomic density of atoms/cm3. This method of conversion relies on a simplified scattering assumption which estimates from the number of collisions expected within a scattering cross section. We take to be the sum of the concentrations of O and N, ignoring contributions from other impurities. This method provides a comparative estimate of across treatments. For reference, the theoretical BCS curves for various were calculated using SRIMP with fixed parameters: K, GHz, nm, and nm Halbritter (1970). Our results are also compared with experimentally determined values from Martinello et al. (2016) and Bafia (2020), obtained through low-energy muon spin rotation (LE-µSR) measurements and cavity fit and extracted with SRIMP Halbritter (1970); Martinello et al. (2016). We observe good agreement between extracted for EP and N doped cavities across three independent measurement techniques: ToF-SIMS, LE-µSR, and cavity fits with RF measurements using SRIMP.
Overall, the data at 16 MV/m ((Fig. 4(b)) highlight two distinct regimes. Treatments that confine impurities to the near-surface, such as baking at and N infusion, primarily reduce without significantly altering the superconducting gap. This results in significant improvement in Eacc but only limited improvement in Q0. It is likely that these treatments form a dirty—clean superconducting bilayer, which can modify the local supercurrent density to act as an additional barrier against vortex penetration Gurevich (2017); Kubo (2016); Lechner et al. (2024). In contrast, N doping reduces and simultaneously increases , resulting in a high Q0 performance. Similarly, longer durations of baking at progressively lower RT by enhancing , even when the average impurity concentration changes minimally. This could be attributed to a homogenization of as observed in point-contact tunneling spectroscopy (PCTS) studies or to changes in the effective penetration depth Groll et al. (2018); McFadden et al. (2023). These results demonstrate that both O and N reduce RT through a combination of mean free path optimization and gap enhancement, with gap enhancement observed only in treatments exhibiting sufficient uniformity in the impurity depth profile throughout the RF layer. A more uniform impurity profile likely ensures that the entire RF penetration layer remains within the optimal scattering regime and enables a more homogeneous superconducting state.
The observed gap enhancements are significantly less pronounced at 5 MV/m, indicating a field dependence consistent with nonequilibrium superconducting effects. Recent studies have shown that in fully gapped superconductors such as niobium, quasiparticle relaxation is dominated by electron–phonon scattering, which reduces the minimum frequency required for nonequilibrium behavior from the 15 GHz predicted by Eliashberg theory to 1.2 GHz Eliashberg (1970); Dobrovolskiy et al. (2020). The cavities investigated in this work operate at 1.3 GHz, placing them within this nonequilibrium regime. This interpretation is further supported by the reported frequency dependence of the anti-Q slope in N doped cavities Martinello et al. (2018). A corresponding systematic study of frequency dependence in O doped cavities would be required to determine whether the observed reduction in RT can likewise be attributed to nonequilibrium superconductivity.
With regard to the hydrogen trapping hypothesis, we observe that N doping increases the average superconducting gap at approximately ten times lower concentration than oxygen. This suggests that N has a stronger effect per atom, potentially due to more effective hydrogen trapping. N has been shown to suppress the formation of niobium nanohydrides Sung et al. (2024). Coincident with this, other studies have demonstrated that N doping reduces the smearing of the quasiparticle density of states compared to EP Bafia et al. (2025b); Groll et al. (2018). These observations linking H minimization and reduced smearing are consistent with theoretical predictions Kubo (2022); Herman and Hlubina (2021); Zarea et al. (2023).
However, first-principle calculations by Ford et al. (2013a) indicate only a minor difference in binding energies for H-O (7.02 eV) and H-N (7.39 eV), which insufficient to explained the observed order-of-magnitude difference. One possible explanation is that N doped surfaces form fewer vacancies and less defective suboxides than O doped surfaces, thereby enhancing their superconducting properties Yang et al. (2018); Wenskat et al. (2022). X-ray photoelectron spectroscopy and transmission electron microscopy studies have also indicated that N doped surfaces form a more stoichiometric Nb2O5 layer with thinner metallic suboxide layers Lechner et al. (2020); Fang et al. (2023). Additional mechanisms may include the higher concentration of interstitial O inducing greater elastic strain on the niobium lattice than the lower concentrations of interstitial N; increases in strain in Nb have been shown to decrease Tc Antoine (2019); Gardner and Smith (1966). Moreover, interstitial oxygen suppresses Tc by K per atomic % ODeSorbo (1963).
In this work, we systematically investigated the effects of O and N impurities on the performance of niobium SRF cavities and their superconducting properties. By correlating material characterizations with RF performance, we demonstrated that O and N independently modify the superconducting properties of the RF layer. N doping achieves comparable RT reduction at lower impurity concentration than O-based treatments. The reduction in RT may be attributable to gap enhancement and homogenization, as the presence of impurities may eliminate defects in the Nb lattice and act as traps for hydrogen. A field dependence in RT is also indicative of nonequilibrium superconductivity, although further studies are necessary to confirm if this is also applicable to oxygen. In comparison, the effect of oxygen is more gradual as it must overcome losses introduced by increased vacancies in the surface oxide. This gradual effect, coupled with the simpler processing requirements for O-based treatments, highlights the potential of oxygen diffusion for finer tuning of the RF surface in future treatments. Additionally, the additive effect of O and N impurities in reducing RT for N infusion indicates potential for combined O and N impurity tailoring. Further theoretical and experimental work is needed to fully understand the distinct roles of oxygen and nitrogen at the atomic scale.
Supplementary Material
See the supplementary material for details on ToF-SIMS analysis, additional ToF-SIMS data on surface contaminants and hydrogen trapping, high-field behavior of RT at 25 MV/m, and a study demonstrating that partial vs fully grown surface oxide has no effect on RT.
ACKNOWLEDGEMENTS
The authors would like to thank Alexandr Netepenko, Adam Clairmont, Akshay Murthy, Damon Bice, Tim Ring, and Davida Smith for their support in sample and cavity preparation, data acquisition, and data analysis. This manuscript has been authored by the FermiForward Discovery Group, LLC under contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work was supported by the University of Chicago.